Calculate the osmotic pressure of a solution containing 16.60 mg of hemoglobin in 15.6 mL of solution at 22 ∘C . The molar mass of hemoglobin is 6.5×104 g/mol.
1
Expert's answer
2020-03-10T09:07:40-0400

"\pi=C_m * R T"

R=0.082atm*litr/mol*K

T=295,15K

Cm="\upsilon" / V [mole/litr]

"\upsilon" =16.6g*10^-3/6,5*10^4 =2.553*10^-7 mole

Cm=1,637*10^-5

"\pi" =0.0003962 atm